OpenStudy (anonymous):

What is the product of five over six and two over three? ten over twelve five over nine seven over nine fifteen over twelve

OpenStudy (anonymous):

write it out.

OpenStudy (anonymous):

hey, umm so sorry to bother again but this is really tricky! @hihi67

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

\[\frac{5}{6} \times \frac{2}{3}= \frac{5 \times 2}{6 \times 3}=?\]

OpenStudy (anonymous):

10/18?

OpenStudy (anonymous):

yes, but you can reduce it.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

10/2=5 18/2=9

OpenStudy (anonymous):

5/9

OpenStudy (anonymous):

YEP!\[\frac{10}{18}=\frac{10 \div 2}{18 \div 2}\]

OpenStudy (anonymous):

BINGO!

OpenStudy (anonymous):

So which one is it?

OpenStudy (anonymous):

OMG! thank you!!!!

OpenStudy (anonymous):

its b

OpenStudy (anonymous):

Yup! Anytime!

OpenStudy (anonymous):

by any chance... are u a teacher?

OpenStudy (anonymous):

no, 9th grader.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Anytime, welcome!

OpenStudy (anonymous):

thx for the medal!

OpenStudy (anonymous):

Oh, it gives me a little fraction of a teamwork score.

OpenStudy (anonymous):

:) see u later im ur fan now

OpenStudy (anonymous):

Ty!